Make New Entry,
Make Followup Entry
User name Markowitz
Log entry time 00:39:09 on July 03, 2010
Entry number 329516
keyword=Swing shift summary
Swing Shift Summary: July 2, 2010
- Pete Markowitz
- Ashot Gasparian
- Start of shift saw the MCC ready to try commissioning of the
corrector coil with the septum magnet. Yves had us ramp up the septum
while the spot on the dump was kept in place by adjusting the corrector
- Septum current of 480 A requires a corrector coil of -4 A
- Septum current of 617.2 A requires a corrector coil of +1 A
- Septum current of 775.6 A requires a corrector coil of +9 A
- The FSD interlocked to the septum current is not working in PSS.
After an access and opening/ closing cabinets, unplugging/ plugging
cables the FSD starts working.
- The FSD interlocked to the corrector coil is not working in PSS.
At end of shift, MCC has called in personnel to debug and fix.